OpenStudy (anonymous):

A linear function passes through the points (-4, 2) and (5, -16). What is the y-intercept?

OpenStudy (anonymous):

f(x) = ax +b f(-4) = 2 f(5) = -16 -4a +b = 2 5a + b = -16 a = -2 b = -6 y-intercept = f(0) = -6

OpenStudy (anonymous):

thx